OpenStudy (kendricklamar2014):

Solve each equation for y. Then find the value of y for each value of x. X - 4y = -4; X = -2, 4, 6

OpenStudy (kendricklamar2014):

@campbell_st @jim_thompson5910

OpenStudy (campbell_st):

well start by adding 4y to both sides of the equation then add 4 to both sides of the equation... what do you get..?

OpenStudy (kendricklamar2014):

I dont get it?

OpenStudy (kendricklamar2014):

@campbell_st

OpenStudy (campbell_st):

ok.... x - 4y = -4 +4y +4y -------------- x = -4 + 4y next add 4 x = - 4 + 4y +4 +4 ---------------- x + 4 = 4y or 4y = x + 4 lastly divide both sides by 4...

OpenStudy (kendricklamar2014):

y = x + 1

OpenStudy (campbell_st):

nearly \[y = \frac{1}{4} x + 1\] now put each x value in to get y start with x = -2 \[y = \frac{1}{4} \times -2 + 1 \] so \[y = \frac{1}{2}\] repeat it for x = 4 and x = 6

OpenStudy (kendricklamar2014):

Is this an X or times ^^^ Y = 1/4 x -2 + 1 ????

OpenStudy (campbell_st):

in this part is times... I've replaced x with -2

OpenStudy (kendricklamar2014):

Y = 1/4 (4) + 1

OpenStudy (kendricklamar2014):

thats how i do it ^^ @campbell_st

OpenStudy (campbell_st):

yes that's it... you should be 2

OpenStudy (kendricklamar2014):

is the answer 2??

OpenStudy (campbell_st):

yes

OpenStudy (kendricklamar2014):

and when u put 6 is the answer 2 1/2

OpenStudy (kendricklamar2014):

@jim_thompson5910 @kirbykirby @HelpBlahBlahBlah

jimthompson5910 (jim_thompson5910):

assuming you mean 6 - 4y = -4 turns into \[\Large y = \frac{5}{2} = 2 \frac{1}{2}\] then you are correct

OpenStudy (kendricklamar2014):

thank u

jimthompson5910 (jim_thompson5910):

np